# Visual Affordances: Enabling Robots to Understand Object Functionality

### v1

@@ III-A Future direction: scaling visual affordance datasets
Available datasets are specific to each affordance redefinition rather than the unified formulation that we provided in this article for visual affordance prediction. Because of this, datasets cannot be easily re-used across different tasks or for the unified case. Moreover, the annotation of object affordances in images and videos is not trivial due to the unclear boundaries of the region on the object, the overlapping of different actions on the same region, and the difficulty of labelling the end-effector pose on objects in the scene. These challenges limit the cross-datasets evaluation of methods and the scalability of datasets for visual affordance, as manual annotations are time-consuming and ambiguous, and requires expensive resources. Because of this, the size of available datasets with curated annotations of visual affordances are limited to less than 50,000 images. As discussed in Section III and Section VI, datasets are used for both training methods and benchmarking their performance. To scale the number of training data, datasets having similar annotation could be merged, adjusting the annotation, or adapting previous methods to provide weakly or self-supervised annotation (e.g. HANDAL [13]). To collect benchmarks, the combination of different methods could help using in-the-wild images with objects in challenging poses and with different backgrounds.
IV Performance measures
In this section, we present the measures to evaluate the performance of models for visual affordance prediction using our framework as a reference. Given the various components of our generic framework, using a single measure is not sufficient to evaluate the performance of a model. Therefore, we discuss a set of performance measures to provide a fine-grained assessment of the various components. For a specific component or sub-task, more than one performance measure can be used to provide a complete assessment and avoid drawing partial or misleading conclusions (see Table VIII). Our discussion refers to performance measures mostly used by previous works, and highlights the characteristics and limitations of these measures (or alternatives). To evaluate our framework, we consider per-class accuracy, per-class precision, per-class recall, and per-class F1 score for functional classification; per-class precision, per-class recall and per-class Jaccard index or Intersection over Union ($IoU$) for functional segmentation; the interpenetration between end-effector and object, and the Analytical grasp score for end-effector pose estimation; and the success rate for experiments with a robot.
@@ IV Performance measures
Functional classification. The performance measures for assessing functional classification are computed across all $N$ samples (image, task, end-effector) of a given dataset, each associated with an affordance category $a$. For each class $a$, a true positive ($TP$) is a sample for which the model predicts the class $a$ and the annotation is also $a$; a false positive ($FP$) is a sample for which the model predicts $a$, but the annotation is a different class; a false negative ($FN$) is a sample with annotation $a$, but the model predicts a different class; a true negative ($TN$) is a sample for which both the model prediction and the annotation are a class different from $a$. Per-class accuracy ($A$) measures the amount of affordance class predictions matching the annotation:
@@
Per-class recall ($R$) measures the amount of affordance class predictions matching the annotation relative to the total number of predictions:
@@
TABLE VIII: Performance measures to evaluate methods for visual affordance prediction. Highlighted in grey the measures we recommend to evaluate the framework.
@@
Accuracy $A$ Eq. 7 $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$
@@
Fréchet Inception Distance $FID$ [133], Eq. 14 $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
@@
KEYS – FUNC: functional classification; FUNS: functional segmentation; EPE: end-effector pose estimation; EIS: end-effector interaction synthesis; ROBV: robot validation; $\bullet$: considered, $\circ$: not considered.
When evaluating the performance of affordance classification methods, previous works [3, 52, 54, 58] showed confusion matrices and computed accuracy. However, the level of detail of confusion matrices makes difficult to quantitatively compare methods. For datasets with imbalanced classes, accuracy is misleading because a high value can be obtained by predicting always the most frequent class. On the contrary, using precision, recall, and F1 provides a complementary analysis while considering imbalanced classes, because precision focuses on false positives and recall on false negatives.
Functional segmentation. The performance measures for assessing the functional segmentation of are per-class precision ($P$), per-class recall ($R$) and per-class Jaccard index or Intersection over Union ($IoU$). To compute these measures, the output probability maps of the model $[0,1]^{W\times H}$ are converted into integer values $\{0,1\}^{W\times H}$ for example using a threshold. As for functional classification, true positives ($TP$), false positives ($FP$), and false negatives ($FN$) are defined for each class $a$. Given the model prediction $\hat{S}$ and the segmentation annotation of the image $S$, a true positive is a pixel $\boldsymbol{y}\in I_{n}$ that is predicted as 1 in $\hat{S}_{n}$ and the corresponding pixel in $S_{n}$ is annotated as 1; a false positive is a pixel $\boldsymbol{y}\in I_{n}$ that is predicted as 1 in $\hat{S}_{n}$ but annotated as 0 in $S_{n}$; a false negative is a pixel $\boldsymbol{y}\in I_{n}$ that is predicted as 0 in $\hat{S}_{n}$, but the corresponding pixel in $S_{n}$ is annotated as 1. Per-class precision measures the percentage of true positives among all positive predicted pixels,
@@
Per-class Jaccard index measures the overlap between predicted and annotated segmentation masks, and quantify how much they are similar in size,
@@
Note that the Jaccard index combines precision and recall, summarising their information in a single value. Therefore, we recommend to report the Jaccard index with complementary performance scores, such as precision and recall, to provide a more comprehensive evaluation and insights.
@@
Most of affordance detection and segmentation works [10, 12, 16, 17, 62, 83, 78, 81, 84, 74, 60] evaluated the performance of methods by using the weighted F-score ($F^{w}_{\beta}$[128]. Weighted F-score weighs false positives based on the Euclidean distance to the closest annotated pixels. However, $F^{w}_{\beta}$ cannot be computed for the classes that are not in the annotated mask, ignoring part of the prediction mistakes. To compare the predicted probability map with the annotation, affordance grounding works [11, 90, 93, 95] used Kullback–Leibler divergence [129], Similarity [130], Normalized Scanpath Saliency [131]. The Kullback-Leibler Divergence has not an upper bound and gives more importance to false negatives compared to false positives. In particular, a false positive results in a Kullback-Leibler Divergence value close to 0, whereas a false negative can cause the value to be high (potentially infinite). Similarity combines together the information of false positives and false negatives, assigning a low value to both errors and hence resulting in an ambiguous interpretation. Normalized Scanpath Saliency considers the prediction values around a neighbourhood of the annotated points. This measure can lead to misleading insights, since the false positives outside the annotation neighbourhood are discarded.
@@
End-effector pose estimation and synthesis. A model can predict an end-effector pose that is plausible and allows a robot to complete the task even if the predicted pose is not annotated. This aspect makes the evaluation of estimated and synthesised end-effector poses challenging. We therefore recommend using interpenetration and analytical grasp score to evaluate the estimated pose, and Fréchet Inception Distance to evaluate the synthesised pose.
The interpenetration is the volume in common between object and end-effector voxels representation (the lower the better) [14]. Note that this performance measure does not consider if the pose of the hand is wrong. The measure cannot be computed if the datasets lacks the annotation of the object pose or the annotation of the end-effector pose. Analytical grasp score [132] computes an approximation of the minimum force to be applied to break the grasp stability by solving a quadratic program. The minimum force corresponds to the smallest Euclidean distance from the origin to any point inside the convex hull composed by all feasible forces and torques combinations. To evaluate the pose of the end-effector (hand), Corona et al. [14] also used the average number of contact fingers: the higher the number of fingers in contact, the stronger the grasp. This measure, however, can penalise actions or objects for which the number of contact fingers is low (e.g. when grasping a glass from the stem). Fréchet Inception Distance ($FID$[133] quantifies the similarity between two Gaussian distributions, one fitted on the synthesised images $\hat{G}\sim(\hat{\mu},\hat{C})$ (where $\mu$ is the mean and $\hat{C}$ the covariance) and the other on the testing set images $G\sim(\mu,C)$ (or ground truth). In particular, the two Gaussian distributions are fitted on the feature representations using the Inception network [134]. $FID$ is computed as:
$FID=||\mu-\hat{\mu}||^{2}_{2}+Tr(C+\hat{C}-2(C\hat{C})^{\frac{1}{2}})\>,$ (14)
@@
where $Tr$ is the trace operator (i.e. the sum of the diagonal elements of a matrix). The first term, $||\mu-\hat{\mu}||^{2}_{2}$, measures the squared difference between the means of the real and generated distributions. A smaller difference indicates that the generated and real images have similar overall features. The second term, $Tr(C+\hat{C}-2(C\hat{C})^{\frac{1}{2}})$, compares the covariances of the real and generated distributions (diversity). A low $FID$ score implies high similarity between the generated images distribution and the testing ones. A high $FID$ score suggests that the distribution of the generated images differs from the distribution of the testing images, either in terms of overall features (mean) or diversity of features (covariance). To evaluate AffordanceDiffusion, Ye et al. [5] also compute contact recall that is the amount of generated hands classified as “in-contact” with the object in the image by an off-the-shelf method [135]. However, in case of unseen objects or unseen conditions (illumination, colour of the background), the method could misclassify whether the hands are in contact or not, leading to a mistake in the computation of contact recall.
@@
Overall evaluation. If a robot is available, the whole framework can be tested in real conditions to assess model performance and identify limitations. In this case, the success rate can be used to measure the performance of the system [10, 83, 81] Reproducing experiments based on success rate is very difficult for some tasks and requires a rigorous protocol. For example, the setup should include information on the object instances, robot model, software versions, and relative poses between object and robot for the task of object lifting. The evaluation should consider separately if the grasping and the lifting are successful, also waiting a fixed amount of time to check if the object falls. When the task is part of other benchmarks [18, 136], using the available performance measures enriches the evaluation.
V Physically-based framework for robot applications
In this section, we present a generic vision-based framework that considers the connection between the affordances and the object properties for robotic manipulation tasks [8]. The framework uses mass estimation as an example of object property. We describe the relationship between affordance and mass and how the mass influences the end-effector interaction with the object. We then discuss the results of previous methods that estimate the mass of a manipulated container regardless of the content.
Figure 9: Visual affordance prediction framework [8]. The vision system predicts the location, the mass and the functional regions of objects in the field of view; based on the predicted information, the control system guides the robotic hand to interact with objects.
V-A Affordance as a function of the object physical properties
To perform the interaction with an object, the robot needs to predict the end-effector pose based on the local geometry of both the end-effector and the object. This requirement for common manipulation tasks, however, does not consider all the physical properties of an object and therefore the action can be successful only in limited conditions. Geometry alone might be insufficient for a robot to estimate a physically plausible grasp pose to complete the task. Objects can vary in their mass [137, 138] because of their material or, in other scenarios, because of a manipulation that can affect the mass. For example, household or kitchen-like objects, such as food boxes, drinking cups and glasses, bowls, bottles, vases, or cleaning products, can be filled with other solid or liquid contents. These objects require the robot to predict a grasp pose that should consider potential unwanted behaviours during the manipulation. For example, the predicted grasp pose should avoid the robot spilling the liquid out of a drinking glass. In a collaborative scenarios, the predicted pose is further constrained by the interaction that another robot or person is performing with the object. Therefore, the range of candidate poses that are commonly estimated in static scenarios (common objects placed on a table top) are not all feasible when other physical properties are also considered and in human-robot collaborations [137, 138]. Affordance prediction depending on relevant object physical properties can affect both motion planning and force regulation and, consequently, influence the interaction between a robot and the object. The local object properties such as shape, stiffness, and material composition, influence the magnitude of the friction force between the hand and the object. The region where to apply the grasping depends not only on the local property of friction but also on the global property of the object weight distribution. The closer is the high friction area to the center of mass, the more stable the grasping will be and the lower the chances that after closing the fingers and lifting the arm, the object will fall [139].
A possibility to consider the mass of the object in the pipeline is to predict the weight from visual data. In the literature, the mass of a filled container is defined as the sum of the mass of the content (filling mass) and the mass of the empty container [137, 138]. The two sub-problems are tackled separately with some methods focusing on the filling mass estimation [140, 141, 142] and other methods focusing on estimating the mass of the empty container [120, 143, 144]. The challenges in estimating the mass of a container range from locating the object in the scene to accurately predict the physical property using only visual data, due to changes of object appearance during manipulation. Manipulated objects may be occluded by the human hand, be altered in shape if made of deformable materials, or be subject to colour change in case the container is transparent and is filled with opaque content.
Fig. 9 shows a framework that relates the mass and the segmentation of affordance regions [8]. Detection methods localise the objects to crop the image and use the object crops as input to other specialised models. The object of interest is selected based on the specific purpose of the interaction, e.g. taking the object that is held by a human. A model specialised in mass estimation predicts the weight of the object and a model specialised in affordance segmentation predicts the regions of interaction. The two independent predictions can be fused to support the movement planning and the adjustment of the robotic hand pose and force during the interaction (Control). Currently, the affordance segmentation and mass estimation are tackled separately due to the unavailability of a single unifying dataset with the appropriate annotation. We analyze the performance of methods estimating the mass to understand what is the current status and what are the main challenges related to the problem of integrating the mass estimation in an affordance prediction pipeline.
V-B Example: estimating the mass of an object
We report the per-category performance of different methods predicting the mass of the container regardless of the content on the private and public testing sets of the CORSMAL Containers Manipulation (CCM) dataset [15, 137, 138]: random sampling (M1), average mass (M2), custom neural network (M3[143], MobileNetV2 with Coordinate Attention (M4[144], and a Neural Network fusing the image features with object aspect ratio and average depth (M5[120]. M1 uses a pseudo-random generator that samples the mass prediction from a uniform distribution in the interval $[1,351]$ using the Mersenne Twister algorithm [145]; M2 computes the average of the mass annotations; M3 [143] predicts the container mass using a custom Neural Network that combines unoccluded segmentation mask, obtained by symmetrically restoring the object mask, with geometric information; M4 uses a MobileNetV2 [146] with Coordinate attention [147] to learn the mass from the object crops (extracted using YOVOv5555https://github.com/ultralytics/yolov5) through an augmentation strategy to vary the dimension of containers based on the mass and to minimize the mass variance for each object category; M5 averages the mass from 5 candidate objects, detected by Mask R-CNN and selected based on the lowest average depth respect to the fixed frontal camera, using a linear layer that processes the concatenation of features extracted from the RGB crops, the aspect ratios of the crops and the average depth.
As performance measures, we use the relative absolute error ($\epsilon$) and the mass estimation score ($s$) between the estimated mass ($\hat{m}^{j}$), and the true mass($m^{j}$[138]. Given a set of recordings $\{v_{j}|j=1,...,V\}$, we compute the relative absolute error as:
${}\epsilon(\hat{m}^{j},m^{j})=\frac{|\hat{m}^{j}-m^{j}|}{m^{j}}\>,$ (15)
were $j$ is the index of a single recording. $\epsilon\in[0,+\infty)$ has a value over $1$ when the estimated mass is greater than the annotated mass, and close to 00 when the estimated mass is lower than the annotated mass. The score $s\in[0,1]$ is used to average the relative absolute error across all $V$ recordings:
$s=\frac{1}{V}\sum_{j=1}^{V}\mathbbm{1}_{j}\>e^{-\epsilon(\hat{m}^{j},m^{j})}.$ (16)
The value of the indicator function $\mathbbm{1}_{j}\in\{0,1\}$ is 0 only when $\hat{m}$ in recording $j$ is not estimated. The score has value 1 when the estimation error is 0 (the predicted mass and annotation are equal).
0020202020404040406060606080808080Score $s$ (%)CCM public testing set
CCM private testing set
cupglassbox00222244446666888810101010Mean $\epsilon$
cupglassbox
Figure 10: Comparison of per-container-type mass estimation score ($s$) and mean of relative absolute error ($\epsilon$) between random sampling ( ), average mass ( ), custom neural network [143] ( ), MobileNetV2 with Coordinate Attention [144] ( ), and neural network fusing image features with object aspect ratio and average depth [120] ( ) on the public and private testing sets of the CORSMAL Containers Manipulation dataset [15, 137]. Note the maximum y-axis value limited to $80$ and $10$, respectively, for visualization purpose. Random obtains $21.43$ and $17.45$ as mean $\epsilon$ for the class cup in the public and private testing sets, respectively, and the average method obtains $12.25$ in the private testing set.
Fig. 10 compares the mass estimation score and the mean relative absolute error $\epsilon$ of the methods, grouped for each container type, on the public and private testing sets of the CCM dataset. On average, models obtain a mass estimation score lower than 56% in the public testing set and 63% in the public testing set (both obtained by M3). This results shows that the models cannot generalise to container instances different from the ones in the training set. For both the public and private testing sets, M1 and M2 achieve higher score values than deep learning based methods for the class glass; in particular, M1 score exceeds other scores by 15 percentage points (p.p.). This result is due to the similar mass of different drinking glasses in the training and testing sets. For the class cup, M5 achieves higher score values than M4 with a performance increase of 24 p.p. in the public testing set and 20 p.p. in the private testing set. In the private testing set, M3 achieves the highest generalization to unseen container instances, outperforming deep learning methods by more than 20 p.p. in the classes cup and box, and more than 2 p.p. in the class glass. Compared to other methods, M3 uses the attention mechanisms and variance minimization to ensure consistent predictions of the same container at test time.
Figure 11: Qualitative results of mass estimation predicted by the method that averages the mass predicted from 5 candidate objects (M5[120]. The first three rows show the RGB, depth and semantic segmentation crops of the objects having the lowest average depth from fixed frontal camera. The prediction is the average of the mass values predicted for the single crops.
Fig. 11 shows M5 mass estimations on 4 different recordings of the CCM testing sets. In case of transparent container (1st column and 3rd column), the detection of the container might fail and also the prediction of the object mask, affecting the mass estimation. In some cases, the selection of candidate containers wrongly selects other objects in the scene, e.g. the pitcher used to fill the containers. These results show that predicting the mass from visual data is a challenging task in case of manipulated containers due to the changes of object appearance and the presence of hand occlusions. The generalization to different object instances in case of occlusions is a key concept that affects also the perception of affordances. An inaccurate estimate of the mass and affordance of an object could lead to a wrong interaction between the end-effector and the object.
V-C Future direction: multimodal estimation of object physical properties
Relating affordance prediction and estimation of object physical properties is far from easy. For example, since humans interact differently based on the objects mass [148], the affordance-mass relation can be learned from human demonstration. Humans have different ways of grasping objects depending on the action they want to perform and how the mass influences the action this section, we showed that estimating the object mass from a visual input is a challenging task and methods cannot yet generalise to unseen objects. Moreover, predicting the mass of a container or other physical properties only from images as input might not be sufficient and other modalities, such as language, audio, and haptic, can be included in our proposed framework [137, 138, 141, 143, 149]. Multimodal models have shown to better generalise to novel and different object categories [44] in tasks such as open-vocabulary object detection [150, 151, 152] and segmentation [153, 154, 155]. For visual affordance, we reviewed and discussed vision-language models addressing mainly affordance grounding and task-driven object detection. The information predicted by these models is coarse but can be used to guide the interaction between hand and object. Audio could complement the visual modality when the appearance of the object is not reliable, e.g. an opaque container whose content is not visible. Haptic could provide a feedback on the local geometry of the object and the force that the end-effector applies on the object to refine the interaction phase [139, 156].
VI Affordance prediction reproducibility
This section discusses the evaluation of affordance predictions and focuses on the reproducibility issues of current benchmarks, resulting in unfair and inconsistent comparisons. Reproducibility is the principle of obtaining the same results given the same conditions (i.e., data, training and testing setups, and trained model) [157, 158]. For models based on neural networks, such as CNN- and Transformer-based models discussed in this article, we also consider the reproducibility of their architecture and learned weights (or model parameters). Reproducibility allows fair comparisons across methods and helps building upon previous works while understanding their limitations. We then highlight open science practices for fair benchmarking.
VI-A Reproducibility challenges
  1. 1.
    data availability for benchmarking (RC1);
  2. 2.
    availability of a method’s implementation (RC2);
  3. 3.
    availability of the trained model (RC3);
  4. 4.
    details of the experimental setups (RC4); and
  5. 5.
    details of the performance measures used for the evaluation (RC5).
In visual affordance prediction, no dataset is collected exclusively for benchmarking methods under specific conditions, such as illumination, clutter, or hand-occlusion (RC1). The majority of previous works [9, 13, 20, 62, 116] trains methods on a training split of data and compares their performance using the testing split of one or more datasets. Cross-dataset evaluations are mostly avoided due to partial overlapping of affordance classes or of object categories, across the selected datasets [4]. As a consequence, researchers train multiple version of the same model, adapted to the classes of a specific dataset. For example, datasets such as UMD [9], IIT-AFF [12], and Multi-View [126] share some of the object and affordance classes, but labelled with different conventions. For example in UMD the class graspable is associated to the class number “1”, while in IIT-AFF to the class number “5”. Additional documentation, such as metadata, help researchers train or evaluate methods only on common categories by re-ordering them. Moreover, relying only on a single benchmark can lead to limited and not generalisable considerations on model rankings. For example, images in UMD and Multi-View are collected in a laboratory environment with static conditions, such as a fixed camera oriented towards a table where an object is placed (camera-object distance is almost always the same) [9, 116]. On the contrary, in real scenarios the camera might be closer or farther from the object of interest compared to the training setting, hence the performance on the benchmark might not reflect the performance on an actual robot.
The lack of publicly available implementation of methods (RC2) [83, 78, 84, 74], the lack of publicly available of trained models (RC3) [12, 62, 83, 78, 81, 84, 74], and the lack of details of experimental setups (RC4) [10, 62, 83, 78, 84, 74] can challenge researchers in reproducing previous works for comparative evaluations. The release of the model trained weights, of the method and inference pipeline implementation, is a crucial aspect for reproducibility, especially for deep-learning based models, allowing other researchers to test the model on their own data without the need for re-training. The availability of model weights is particularly important when researchers need a comparison, as re-training the model can be too time- and resource-consuming. In case a new dataset is proposed and a previous method needs re-training, only the method implementation is sufficient, while the trained weights are not necessary. The re-implementation of methods and setup can often be time-consuming and prone to errors, and not always resulting in the expected outcome (i.e., results are not replicable or findings are not reproducible). To avoid this issue and often to save time, researchers choose to report the results from previous works as they are [83, 78, 84, 74]. This practice can result in an unfair comparison if the experimental conditions are not the same, leading to misleading findings and conclusions.
TABLE IX: Comparison of training/testing setups used by different methods for affordance detection and segmentation on the UMD dataset [9] along with their performance [28]. We selected the backbone providing the highest $F^{w}_{\beta}$. Due to the setup inconsistencies, the straight comparison among models performance is unfair.
Training setup Resolution Data augmentation Image resize procedure $F^{w}_{\beta}$
Flipping Scaling Rotating Jittering Training set Testing set
AffordanceNet [10] $1000\times 600$circ $\circ$ $\circ$ $\circ$ unknown unknown 79.90
CNN [62] $320\times 240$circ $\circ$ $\circ$ $\circ$ centre-crop sliding window 76.60
DRNAtt [78] $320\times 240$circ $\circ$ $\circ$ $\circ$ centre-crop unknown 94.10
RANet [84] $224\times 224$circ $\circ$ $\circ$ $\circ$ centre-crop unknown 86.13
GSE [74] $400\times 400$bullet $\bullet$ $\circ$ $\circ$ crop unknown 85.50
BPN [83] $1000\times 600$bullet $\bullet$ $\bullet$ $\bullet$ unknown unknown 86.21
KEYS – $\bullet$: considered, $\circ$: not considered, Jittering: colour jittering
Using the same experimental setup to train and test affordance models allows a fair comparison whose analysis can better validate the technical contributions proposed by a novel work. When releasing the training and testing code is not possible, reporting all the training and testing details to reproduce a setup becomes fundamental, enabling other researchers to reimplement the method and correctly compare their solution. The experimental setup details include training hyper-parameter values, the chosen data splits, image pre-processing (normalisation and cropping procedures) and post-processing. The lack of details of the experimental setup causes methods for affordance detection and segmentation to be often not reproducible [10, 62, 83, 78, 84, 74]. For example, AffordanceNet and BPN do not include the detail of the input image resize during training and testing phases [10, 83], whereas DRNAtt, RANet, and GSE do not include these details during the testing phase. Other details often omitted are the parameters of the optimizers used during training [78, 90, 93, 94]. Apicella et al.’s work [4] showed that these lack of details in the experimental setup led to unfair and inconsistent comparisons.
Previous works evaluate the performance of different methods using scores or metrics that quantify the discrepancy between predictions and annotations. Describing a performance measure help other researchers understand if the experiment validates their claim or if a different measure should be chosen. Providing the mathematical formulation of the scores helps disambiguate similar meaning but different implementations, especially when a public evaluation toolkit is not used or referred to. For example, mean Intersection over Union (IoU) can be the average of all the IoUs between prediction and annotation, or the IoU considering the full set of predictions and annotations. Previous works evaluated a few methods with different performance measures, making comparison and ranking not possible. For example, AdaptiveNet [159] and STRAP [86] compared their performance with baselines on CAD120-AFF using IoU as a performance measure. However, these works excluded the comparison with other available models that were mostly tested on UMD but using $F^{w}_{\beta}$ as a performance measure.
Affordance detection and segmentation methods are difficult to reproduce due to missing implementation and lack of setups details [10, 62, 83, 78, 84, 74]. We report the training and testing setups of affordance detection and segmentation methods on the UMD dataset in Table IX, along with the average performance from Chen et al.’s survey [28]. Despite being trained and tested on the same dataset, models’ performance is not directly comparable due to inconsistencies in the setups such as the image resize procedure (image cropping or input resolution) and augmentation procedure during training.
Inconsistencies can also be present previous method tackling a similar problem are adapted into a baseline to compare with. For example, because of the missing annotation of the object pose in the training set, AffordanceDiffusion [5] is compared with the coarse hand prediction of GanHand [14]. Despite providing implementation details as supplementary material, the results suggest that the whole GanHand was used for comparison. However, since a part of the architecture and of the training procedure is missing, this adapted version of GanHand is only a proxy to the actual (unknown) performance of GanHand.

### v2

@@ VI Future directions
Scaling visual affordance datasets. Datasets cannot be easily re-used across different tasks or for the unified case, as each dataset is specific to an affordance redefinition rather than the unified formulation. Moreover, the annotation of object affordances in images and videos is not trivial due to the unclear boundaries of the region on the object, the overlapping of different actions on the same region, and the difficulty of labelling the agent’s hand pose on objects in the scene. These challenges limit the cross-datasets evaluation of methods and the scalability of datasets for visual affordance, as manual annotations are time-consuming and ambiguous, and require expensive resources (as discussed in Sec. IV and Sec. V). To scale the number of training data, datasets having similar annotation could be merged, adjusting the annotation, or adapting previous methods to provide weakly or self-supervised annotation (e.g. HANDAL [7]). The combination of different methods could help using in-the-wild images with objects in challenging poses and with different backgrounds.
Benchmarking visual affordance. Reproducibility and advancements in the design of novel solutions has been facilitated by available datasets, benchmarks and competitions in various computer vision tasks (e.g. BOP for object 6D pose estimation [141]). However, benchmarks for visual affordance predictions are not yet available. Nevertheless, solutions based on our generic formulation and novel methods can be designed for robotic grasping and manipulation tasks [142], picking in clutter [143], and human-to-robot object handovers [123], whose benchmarking protocols and competitions are available. A benchmarking protocol specific to visual affordance could be designed and included in existing competitions to further promote reproducibility and engagement.
Appendix A Performance measures
We detail the measures to evaluate the performance of models for visual affordance prediction using our formulation as a reference. For each component or sub-task, more than one performance measure can provide a complete assessment and avoid drawing partial or misleading conclusions (see Table X). We highlight the characteristics and limitations of performance measures used by previous works.
@@ Appendix A Performance measures
Functional classification. The performance measures for assessing functional classification are computed across all $N$ samples (image, task, hand) of a given dataset, each associated with an affordance category $a$. For each category, a true positive ($TP$) is a sample for which the model predicts $a$ and the annotation is also $a$; a false positive ($FP$) is a sample for which the model predicts $a$, but the annotation is a different class; a false negative ($FN$) is a sample with annotation $a$, but the model predicts a different class; a true negative ($TN$) is a sample for which both the model prediction and the annotation are a class different from $a$. Per-class accuracy ($A$) measures the amount of affordance predictions matching the annotations:
@@
Per-class recall ($R$) measures the amount of class predictions matching the annotations among all class annotations:
@@
TABLE X: Performance measures to evaluate methods for visual affordance prediction. Highlighted in grey the measures we recommend for evaluation.
@@
Accuracy AA Eq. 7 $\bullet$ $\circ$ $\circ$ $\circ$ $\circ$
@@
Fréchet Inception Distance FIDFID [149], Eq. 14 $\circ$ $\circ$ $\circ$ $\bullet$ $\circ$
@@
KEYS – FUNC: functional classification; FUNS: functional segmentation; EPE: hand pose estimation; EIS: hand interaction synthesis; ROBV: robot validation; $\bullet$: considered, $\circ$: not considered.
When evaluating the performance of affordance classification methods, previous works [3, 44, 46, 50] showed confusion matrices and accuracy. However, the level of detail of confusion matrices makes it difficult to quantitatively compare methods. For datasets with imbalanced classes, accuracy is misleading because a high value can be obtained by predicting always the most frequent class. On the contrary, using precision, recall, and F1 provides a complementary analysis while considering imbalanced classes, because precision focuses on false positives and recall on false negatives.
Functional segmentation. The performance measures for assessing the functional segmentation of are per-class precision ($P$), per-class recall ($R$) and per-class Jaccard index ($J$) or Intersection over Union ($IoU$). To compute these measures, the output probability maps of the model $0,1]^{W\times H}$ are converted into integer values $\{0,1\}^{W\times H}$ for example using a threshold. As for functional classification, true positives ($TP$), false positives ($FP$), and false negatives ($FN$) are defined for each class $a$. Given the model prediction $\hat{S}$ and the segmentation annotation of the image $S$, a true positive is a pixel $\boldsymbol{y}\in I_{n}$ that is predicted as 1 in $\hat{S}_{n}$ and the corresponding pixel in $S_{n}$ is annotated as 1; a false positive is a pixel $\boldsymbol{y}\in I_{n}$ that is predicted as 1 in $\hat{S}_{n}$ but annotated as 0 in $S_{n}$; a false negative is a pixel $\boldsymbol{y}\in I_{n}$ that is predicted as 0 in $\hat{S}_{n}$, but the corresponding pixel in $S_{n}$ is annotated as 1. Per-class precision measures the percentage of true positives among all positive predicted pixels,
@@
Per-class Jaccard index combines precision and recall measuring the overlap between predicted and annotated segmentation masks, and quantifying how much they are similar in size,
@@
We recommend to report the Jaccard index with complementary performance scores, such as precision and recall, to provide a more comprehensive evaluation and insights.
@@
Most affordance detection and segmentation works [12, 6, 10, 56, 54, 77, 72, 75, 78, 68, 52] evaluated the performance of methods using the weighted F-score ($F^{w}_{\beta}$[144]. $F^{w}_{\beta}$ weighs false positives based on the Euclidean distance to the closest annotated pixels, ignoring the classes that are not in the annotated mask. To compare the predicted probability map with the annotation, affordance grounding works [13, 84, 88, 90] used Kullback–Leibler divergence [145], Similarity [146], Normalized Scanpath Saliency [147]. The Kullback-Leibler Divergence gives more importance to false negatives compared to false positives. In particular, a false positive results in a Kullback-Leibler Divergence value close to 0, whereas a false negative can cause the value to be high (potentially infinite). Similarity combines together the information of false positives and false negatives, assigning a low value to both errors and hence resulting in an ambiguous interpretation. Normalized Scanpath Saliency considers the prediction values around a neighbourhood of the annotated points. This measure can lead to misleading insights, since the false positives outside the annotation neighbourhood are discarded.
@@
Hand pose estimation and synthesis. Predicted poses of the hand, also different from the annotated ones, can enable a robot to complete the task, making the evaluation of estimated and synthesised hand poses challenging. We therefore recommend using interpenetration and analytical grasp score to evaluate the estimated pose, and Fréchet Inception Distance to evaluate the synthesised pose.
[…]
@@
where $Tr$ is the trace operator (i.e. the sum of the diagonal elements of a matrix). The first term, $\mu-\hat{\mu}||^{2}_{2}$, measures the squared difference between the means of the real and generated distributions. A smaller difference indicates that the generated and real images have similar overall features. The second term, $Tr(C+\hat{C}-2(C\hat{C})^{\frac{1}{2}})$, compares the covariances of the real and generated distributions (diversity). A low $FID$ score implies high similarity between the generated images distribution and the testing ones. A high $FID$ score suggests that the distribution of the generated images differs from the distribution of the testing images, either in terms of overall features (mean) or diversity of features (covariance). To evaluate AffordanceDiffusion, Ye et al. [5] also compute contact recall that is the amount of generated hands classified as “in-contact” with the object in the image by an off-the-shelf method [151]. However, in case of unseen objects or unseen conditions (illumination, colour of the background), the method could misclassify whether the hands are in contact or not, leading to a mistake in the computation of contact recall.
@@
Overall evaluation. If a robot is available, models performance can be assessed in real conditions using success rate [12, 75, 77]. Reproducing experiments based on success rate is difficult for some tasks and requires a rigorous protocol. The setup should include information on the object instances, robot model, software versions, and relative poses between object and robot. The evaluation should consider separately if actions are successful (e.g., grasping and lifting), also waiting a fixed amount of time to check if the object falls. When the task is part of other benchmarks [123], using the available performance measures enriches the evaluation.
Appendix B Affordance sheets
We provide some examples of compiled affordance sheets for related works based on the available information [38, 43, 100]. In particular, VLTP [38] is a method for object localisation, EgoTopo [43] for functional classification, and Multi-FinGAN [100] for Hand pose estimation.